# Fault Diagnosis and Prognosis Capabilities for Wind Turbine Hydraulic Pitch Systems: An Overview

### v1

@@ 5 Discussion
In terms of isolability, two groups of violations of constraints were shown to be group-wise isolable. The first group was represented by defects related to the valve, i.e., changes in voltage or friction (Equation 13). The second referred to leakages on the supply line, i.e., rotary unit leakage, relief valve defect, or accumulator defect. All other defects that were inserted into the system were shown to be isolable, which means that standard residual generators will suffice to detect and isolate the root cause of the changes they create in the pitch system signals.
@@
The outcome of the detectability analysis was that the only fault that couldn’t be detected from a structural perspective is an increased friction in the cylinder. This was shown to be due to the presence of a disturbance term in the same equation. This result from the structural analysis is very useful indeed. It can be used to automatically generate residuals that achieve very convincing coverage of detectability and isolability, and it points to the one component in the system where supplemental effort is needed for condition monitoring. The structural result does not imply that changes in cylinder friction could not be detected. It just shows that the residuals auto-generated by structural analysis would not suffice to capture changes in cylinder friction separately from variations in load torques on the turbine blades from wind. The estimation of cylinder friction is essential for pitch system condition monitoring and will be subject for further scrutiny, (dallabonaFrictionEstimationCondition2024, ).
@@
The topology of a system may change in different operational modes, or according to system excitation via different paths of control. This can be utilized to enhance fault isolation over what is visible from the standard structural isolability analysis. Active fault isolation is a set of techniques where control inputs are excited and/or operational mode is changed, to explore different input-to-output paths in topology. Structural results for active fault isolation were presented in (blankeStructuralDesignSystems2007, ), and enhanced isolation by shifting between operational modes, was shown for a water-hydraulic medical ”water for injection” plant in (laursenFaultDiagnosisWater2008a, ). Methods for signal-based active isolation enhance the structural results by using the time/frequency domain properties, which structural analysis does not necessarily show. A generic setup for active fault diagnosis was published in (poulsenActiveFaultDiagnosis2008, ), with demonstration of active diagnosis on a high fidelity wind turbine simulation in (niemannFaultDiagnosisCondition2018, ) where continuous-time analysis and linear system assumptions were utilized.
@@ 6 Conclusions
This article provided an extended model that can be used for fault diagnosis and fault tolerant control in hydraulic pitch systems of wind turbines. It merged the dynamics during normal operation and all the major faults that could occur in the system’s lifetime, together with a description of their development and signal properties. The combination of these two pieces of information was leveraged for deriving the structural detectability and isolability properties achievable with the standard configuration presented, by means of a Dulmage-Mendelson decomposition. Finally, different sensor configurations were tested for reducing the level of redundancy while maintaining the same isolability performances. The outcome was an indication of the minimal sensor requirement for being able to maximize fault structural isolability in the most generic scenario of wind turbine hydraulic pitching.

### v2

@@ 6 Discussion
In terms of isolability, two groups of violations of constraints were shown to be group-wise isolable. The first group was represented by changes in voltage or friction in the valve (Equation 14). The second referred to leakages on the supply line, i.e., rotary unit leakage, relief valve defect, or accumulator defect. All other defects that were inserted into the system were shown to be isolable, which means that standard residual generators will suffice to detect and isolate the root cause of the changes they create in the pitch system signals.
@@
The outcome of the detectability analysis was that the only fault that couldn’t be detected from a structural perspective is an increased friction in the cylinder. This was shown to be due to the presence of a disturbance term in the same equation. This result from the structural analysis can be used to automatically generate residuals that achieve very convincing coverage of detectability and isolability, and it points to the one component in the system where supplemental effort is needed for condition monitoring. The structural result does not imply that changes in cylinder friction could not be detected. It merely shows that the residuals generated by structural analysis will not discriminate changes in the cylinder friction from variations in load torques exerted on the turbine blades from wind. The estimation of cylinder friction is essential for pitch system condition monitoring and has been studied in [14]. Finally, it should be noted that the topology of a system may change in different operational modes, or according to system excitation via different paths of control. This can be utilized to enhance fault isolation over what is visible from the standard structural isolability analysis.
@@
Another approach to improve on the effectiveness of the structure-based diagnosis, pertains to active fault isolation, i.e., a set of techniques where control inputs are excited and/or operational mode is changed, to explore different input-to-output paths in topology. Structural results for active fault isolation were presented in [7], and enhanced isolation by shifting between operational modes, was shown for a water-hydraulic medical ”water for injection” plant in [34]. Methods for signal-based active isolation enhance the structural results by using the time/frequency domain properties, which structural analysis does not necessarily show. A generic setup for active fault diagnosis was published in [53], and tested on a high fidelity wind turbine simulation in [45] where continuous-time analysis and linear system assumptions were utilized.
As discussed in Section 5, the robustness of fault detection can be compromised by parametric and model uncertainty. Condition (28) relies on the calculation of bounds on conservative bounds on the gradients $\frac{\partial\breve{r}}{\partial p}$, $\frac{\partial\breve{r}}{\partial d}$ and $\frac{\partial\breve{r}}{\partial f}$, which could limit what can be detected. Methods based on signal processing and statistical change detection offer supplementary features that enhance diagnosability of faults even in the presence of noise. For example, a test statistics $g_{r}$ of $r(t)$ can be compared with a threshold $h$ that is obtained by imposing requirements on probabilities of detection and false alarms [5]. Additional robustness can be achieved without sacrificing sensitivity by using commissioning data to obtain adaptive thresholds.
@@ 7 Conclusions
This article provided an extended model that can be used for fault diagnosis and fault tolerant control in hydraulic pitch systems of wind turbines. A wide range of major faults that can occur in the system’s lifetime, were integrated into the model that corresponds to dynamics during nomimal operation. This integration comprised information on the faults’ signal properties as well as their change profiles. The combination of these two pieces of information was leveraged for deriving the structural detectability and isolability properties that are achievable with the presented configuration, i.e., by means of a Dulmage-Mendelson decomposition. Different sensor configurations were tested to reduce the level of redundancy while maintaining the same isolability performances. The outcome was an indication of the minimal sensor requirement for facilitating maximum fault structural isolability in the most generic scenario of wind turbine hydraulic pitching. Finally, elements of the robustness of the presented approach with respect to parametric uncertainty were discussed in a test cases. The results highlighted that although the quality of fault detection and estimation is subject to variations in the model parameters, the latter do not affect the structural detectability and isolability features of the system.
Future work will include a full-scale sensitivity analysis to quantify the effects of parametric uncertainty on the diagnostic performance of the generated residuals. Furthermore, active fault diagnosis will be considered in connection to potential improvements on the detection and isolation capabilities of the system. Experimental verification of the results will be pursued by implementing structure-based residual generators on a real pitch hydraulic system.